Comparative Statements of Retained Earnings for Renn-Dever Corporation were reported as follows for the fiscal years ending December 31, 2016, 2017, and 2018.

RENN-DEVER CORPORATION
Statements of Retained Earnings For the Years Ended December 31
2018 2017 2016
Balance at beginning of year $ 6,984,092 $ 5,565,852 $ 5,734,552
Net income (loss) 3,319,700 2,350,900 (168,700 )
Deductions:
Stock dividend (38,000 shares) 253,000
Common shares retired, September 30 (150,000 shares) 223,660
Common stock cash dividends 900,950 709,000 0
Balance at end of year $ 9,149,842 $ 6,984,092 $ 5,565,852


At December 31, 2015, paid-in capital consisted of the following:

Common stock, 2,050,000 shares at $1 par, $ 2,050,000
Paid in capital—excess of par 7,530,000


No preferred stock or potential common shares were outstanding during any of the periods shown.

Required:
Compute Renn-Dever’s earnings per share as it would have appeared in income statements for the years ended December 31, 2016, 2017, and 2018.
Answer #1

Number of shares outstanding :

31/12/2015 2,050,000
31/12/2016 2,050,000
31/12/2017 2,050,000-150,000 = 1,900,000
31/12/2018 1,900,000+38,000 = 1,938,000

Calculation of weighted average number of shares : number of shares * period of outstanding

In case dates are not mentioned we use Average number of shares as weighted average shares =[Beginning shares+ending shares ]/2

weighted average number of shares
2016 [2050000+2050000]/2= 2050000
2017

[2050000*9/12](1Jan -30Sep ) +[1900000*3/12] (1Oct-31Dec)

1537500+ 475000

2012500

2018

[1900000+1938000]/2

1919000

Earning per share is a profitability ratio which reports the amount of profit/(loss) per share attributable to common stockholders .

It is calculated using the following formula :[Net income/(loss) - preferred dividend paid if any ]/weighted average number of shares .

Earning per share
2016

[-168700 -0]/2050000

-168700/2050000

-.08 per share

2017`

[2350900 - 0]/2012500

2350900/2012500

$ 1.17 per share

2018

[3319700-0]/1919000

=$ 1.73 per share